find the answer to make x the subject in a(x+1)=b(x+2)
Answer by nerdybill(7384)

a(x+1)=b(x+2)
ax+a=bx+2b
ax=bx+2b-a
ax-bx=2b-a
x(a-b)=2b-a
x=(2b-a)/(a-b)